# The resummation model in FLAMINGO: precisely predicting matter power suppression from observed halo baryon fractions

### v1

@@ 4.3 Other applications: halo mass function
We show the result of applying the mass-independent, universal $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation of Fig. 4 to the halo masses of several FLAMINGO simulations in Fig. 16, binning the resulting masses to obtain the HMF. We show the HMF down to the mass scale corresponding to $100$ DMO particles, and the three most massive bins are combined into one to ensure that every bin contains at least $10$ DMO haloes. In the top part of the figure we show the HMF itself, and in the bottom part we examine the relative difference between all HMFs and the DMO HMF. Grey shaded regions show deviations of $2\%$ and $5\%$. The results for the unscaled hydro masses are similar to the results presented in Schaye et al. (2023, their Fig. 20), but now presented for $M_{\mathrm{500c}}$: the different FLAMINGO simulations shift the HMF downward by up to $\approx 20\%$ due to feedback (solid lines), with the effect extending to higher halo masses for simulations with stronger AGN feedback. However, when we divide the masses by the predicted retained mass fraction for their mean halo baryon fractions (dashed lines), the effects of feedback are compensated over more than two decades of halo mass, and we recover the DMO HMF to percent-level precision. While the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation used here was fit only to haloes from the L2p8_m9 simulation with $M_{\mathrm{500c}}\geq 10^{12}\,\mathrm{M_{\odot}}$, it applies to all FLAMINGO simulations, and divergence at lower halo masses is slow.
Figure 16: The $M_{\mathrm{500c}}$ halo mass function (HMF) for L1_m9_DMO and various hydrodynamical variations. Shown as dashed lines are the HMFs we obtain when we divide the halo masses in the hydrodynamical simulations by their expected retained mass fractions, obtained from their baryon fractions. The bottom panel shows that without this correction (solid lines), the HMF shifts downward by up to $\approx 20\%$ depending on the strength of feedback, over a wide range of mass scales. Grey shaded regions show deviations of $2\%$ and $5\%$. After correcting the hydrodynamical masses with the retained mass fractions (dashed lines), the DMO HMF is recovered to percent-level precision over more than two decades of mass for all simulations.
We note that Euclid Collaboration: Castro et al. (2024) recently also explored a relation to map halo masses to their DMO equivalent to recover the (cumulative) HMF, and they too start from what we call the corrected baryon fraction. However, their equivalent of the retained mass fraction is defined by adding a parameter $\delta_{f}$ to the measured baryon fraction instead, which they calibrate to simulations. The precision they obtain is similar to ours, however, the relation of the current work is applicable over a much wider range in halo mass scales and feedback strengths, without requiring recalibration.
@@ 4.4 Strengths and limitations
The main strength of the resummation model is that it allows halo baryon fractions to be mapped directly to a matter power suppression signal with zero free parameters, meaning that observed baryon fractions could be straightforwardly marginalised over to give a confidence interval around the predicted suppression. In this way it is similar to SP(k) (Salcido et al., 2023) – however, the resummation model does not assume any scaling law between baryon fractions and halo masses, and is much more flexible in the suppression signals it can generate, having not been trained on any simulated suppression signals.
@@
Compared to the baryonification/baryon-correction models (e.g. Schneider & Teyssier, 2015; Schneider et al., 2019; Aricò et al., 2020, 2021; Schneider et al., 2025), which can also predict quantities beyond the suppression signal and can operate at the field level, the applicability of the resummation model is more narrow. However, the resummation model again has zero free parameters, at least at low redshift, and does not assume any functional forms for density profile shapes, whereas the default baryonification models have seven free parameters and do require these assumptions. For the same reasons, the resummation model compares favourably to halo-model based approaches (e.g. Semboloni et al., 2011; Zentner et al., 2013; Debackere et al., 2020; Pandey et al., 2025), including the popular hmcode (Mead et al., 2015, 2021), which contain many parameters that are trained on a limited set of DMO and hydrodynamical simulations, typically leaving at least one parameter modelling the strength of AGN feedback free. The resummation model, then, offers a great alternative to less flexible models for the suppression signal, and could be complementary to baryonification-based approaches: using its predictions for the retained mass fractions and suppression signals may help constrain the parameter space of baryonification models, allowing for more precise predictions of other quantities like halo pressure and temperature profiles.
@@
The main underlying assumption of our model is that the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation – linking the mean halo baryon fraction to the fraction of mass lost due to baryonic effects, relative to DMO – which applies to all well-resolved FLAMINGO haloes regardless of their baryonic physics, is truly universal, i.e. that it also applies to haloes in the real Universe. We have checked that haloes from the cosmo-OWLS (Le Brun et al., 2014) and BAHAMAS (McCarthy et al., 2017) simulations indeed lie on top of the same relation we present here – despite their cosmic baryon fractions lying outside the range probed by FLAMINGO– and encourage other simulators to do the same. The results of Appendix D do indicate that the relation may shift by $\sim 1$ percentage point for lower-mass haloes when the simulation’s mass resolution is improved by a factor of 8, which we will investigate further in future work.
@@
The second-most important assumption of our model is that the non-resolved-halo matter component is sub-dominant on non-linear scales, and that we therefore do not need to accurately model the distribution of ejected matter. As we have shown in §3.6, this assumption breaks down at high redshifts, although our results there indicate that adding a possibly also universal but redshift-dependent parameter ($f_{\mathrm{ret,5x500c}}$}) could solve this issue. Some less important assumptions taken include that the linear halo bias is preserved as haloes lose mass (though we find that the relative change in bias is typically $\lesssim 1\%$), and that the halo positions do not change due to feedback in a way that significantly impacts their clustering (justified by the results of van Daalen et al., 2014; Zennaro et al., 2024).
@@ 5 Summary
In order to derive unbiased cosmological parameters from Stage-IV weak lensing surveys like Euclid, we need models that can predict the matter power spectrum up to at least $k\,{\approx}\,{10}\,h\mathrm{\,Mpc^{-1}}$ with percent-level precision. The main challenge in this endeavour, as first put forward in van Daalen et al. (2011), is that baryonic feedback significantly redistributes matter on large scales, relative to dark matter alone, the details of which are still not understood to the required level of precision. Many modelling efforts and mitigation strategies based on the results of hydrodynamical simulations exist, but these all assume functional forms for halo density profiles or the suppression signal itself, have free parameters that need to be marginalised over, or both. Here, we presented an improved version of the resummation model first described in van Loon & van Daalen (2024), which maps observed halo baryon fractions of groups and clusters to a flexible suppression signal without any free parameters, with a typical precision of $\lesssim 1\%$.
@@
Conceptually, the model takes observed halo masses and baryon fractions within one or more spherical overdensity radii $R_{\Delta}$, translates these into retained mass fractions $f_{\mathrm{ret},i,\Delta}\equiv M_{\mathrm{hydro},i,\Delta}/M_{\mathrm{DMO},i,\Delta}$ for different halo mass bins $i$ through the mass-independent, feedback-independent $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation presented in §2.2.1, and uses these to rescale DMO halo-matter cross spectra to account for the mass lost due to baryonic feedback. The mass removed from haloes is assumed to cluster similarly to matter already outside haloes, allowing us to rescale the DMO non-resolved-halo cross spectrum as well. All these cross spectra are then resummed to obtain a new matter power spectrum including baryons. Extensions to larger or smaller scales are straightforward (see §2.2.2-2.2.3), and allow for $\sim 1\%$ precision up to $k\,{\approx}\,{25}\,h\mathrm{\,Mpc^{-1}}$ given stellar and gas fractions within the inner halo. The steps of the model are summarised in §2.2.4, and a Python package called resummation will be made available upon acceptance of this work for publication. For a given FLAMINGO cosmology, the model maps halo baryon fractions to power suppression predictions in $1\,\mathrm{ms}$, allowing for fast iterations to marginalise over the uncertainties of observed baryon fractions of each mass bin independently.
@@
We have quantified the contributions to the suppression signal of haloes of different masses (§3.4), explaining and confirming the findings of e.g. van Daalen et al. (2020) on the dominant role played by haloes around $M_{\mathrm{500c}}=10^{14}\,\mathrm{M_{\odot}}$, in line with van Loon & van Daalen (2024). We have shown that the baryon fraction within $R_{\mathrm{500c}}$ is also a strong predictor of the retained mass fractions on larger scales as well (§3.5), allowing the resummation model to yield accurate predictions even when only $f_{\mathrm{b,500c}}(M_{\mathrm{500c}})$ is available. We have also shown that the model’s assumption about the minor role of non-resolved-halo matter starts to break down at high redshift, but that this can be accounted for as well (§3.6) – though whether this too is model-independent remains to be seen. Finally, in Section 4, we have shown that halo relaxation following ejection likely shapes the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation (§4.1), applied our model to observed baryon fractions within $R_{\mathrm{500c}}$ to show the suppression signals they predict (§4.2), shown that the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation can also be used to accurately recover the DMO HMF from the observed halo mass function (§4.3), and discussed the strengths and limitations of the model (§4.4).
@@
Future work on the resummation model could include: improved modelling of the non-resolved-halo component, particularly for high $z$; emulation of DMO halo-matter cross spectra; investigating the resolution dependence of the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation, and its dependence on $\Omega_{\mathrm{b}}/\Omega_{\mathrm{m}}$ outside the limited range probed by FLAMINGO; reformulating the model in terms of multiples of $R_{\mathrm{500c}}$ and/or more readily observable halo mass definitions; inverting the model to yield baryon fractions given a suppression signal, without degeneracies; and improving the model in the presence of asymmetric redistributions of matter, such as through jets.
@@ Acknowledgements
We thank Jaime Salcido for providing the baryon fractions and Monte Carlo errors derived from Akino et al. (2022), converted to the D3A cosmology. We thank Ajay Dev for providing the baryon fractions and percentiles derived from the combined dataset of Dev et al. (2024) and a useful discussion on their cosmological and mass definition dependencies.
@@
The halo-matter DMO cross spectra, as well as baryon fractions and stellar fractions for the FLAMINGO simulations used in this work, will be made publicly available as part of the Python package upon acceptance of this work for publication. The matter power spectra of all FLAMINGO simulations are already available on the FLAMINGO project’s web page. [^14] https://flamingo.strw.leidenuniv.nl/
@@ Appendix C Fitted results at higher redshifts
That there exists a redshift dependence at all, hints at other dependencies of the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation. Elbers et al. (2025) show that the retained mass fraction, there indicated as $R_{\mathrm{b}}$, depends on halo concentration and formation time at fixed DMO mass. While the relation presented in this work likely absorbs some of those dependencies, some residual dependence probably remains.
@@ Appendix D Sensitivity to changes in resolution
The higher-resolution L1_m8 simulation also agrees for massive haloes. However, as the halo mass decreases it starts to undershoot the m9 simulations by $1$ percentage point consistently for $R_{\mathrm{200m}}$ regions. For $R_{\mathrm{500c}}$, the deviation is larger, though it still does not exceed $\approx 2$ percentage points at any halo mass. A similar discrepancy between mass ratios in these FLAMINGO simulations was recently noted by Schneider et al. (2025, their Appendix B).
@@
We compare the model prediction for the suppression signal for L1_m8 to the true one in Fig. 20, using the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation fit to L2p8_m9. The model tends to underestimate the amount of suppression, as expected from the results of Fig. 19, since the retained mass fraction is overestimated. Still, the prediction is $\lesssim 1\%$ precise for $k\,{\lesssim}\,{3}\,h\mathrm{\,Mpc^{-1}}$, and stays below $4\%$ for all $k\,{\leq}\,{10}\,h\mathrm{\,Mpc^{-1}}$.
We note that the resolution is not the only property of the simulation that changes between L1_m9 and L1_m8: as noted in Schaye et al. (2023), two changes were made to the simulation code after most of the FLAMINGO simulations were completed. First, there was an update to the simulation code that fixed the repositioning of supermassive black holes, to keep them in the centre of their halo’s potential. In most FLAMINGO simulations, the black hole’s own potential was erroneously not excluded from the potential calculation (see their §2.3.5). Second, in these same simulations, particles with exactly zero metallicity had their star formation suppressed, although this only significantly affected the very lowest-mass galaxies (see their footnote 5). Of the simulations explored here, only L1_m8, L1_m10 and the jet simulations have the black hole potential correctly subtracted, and the star formation bug fixed. In our results, these are among the most divergent simulations in terms of model performance, which the code changes (particularly the black hole repositioning) could possibly have contributed to.
Figure 20: The result of applying the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation fit to L2p8_m9 to the higher-resolution L1_m8 simulation. Left: The true suppression signal compared to the model prediction. Right: The ratio of the model to true suppression signal. As expected from the results of Fig. 19, the universal relation overpredicts the retained mass fraction for this resolution, although the prediction is still $\lesssim 1\%$ precise for $k\,{\lesssim}\,{3}\,h\mathrm{\,Mpc^{-1}}$.

### v2

@@ 4.3 Other applications: halo mass function
We show the result of applying the mass-independent, “universal” $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation of Fig. 4 to the halo masses of several FLAMINGO simulations in Fig. 17, binning the resulting masses to obtain the HMF. We show the HMF down to the mass scale corresponding to $100$ DMO particles, and the three most massive bins are combined into one to ensure that every bin contains at least $10$ DMO haloes. In the top part of the figure we show the HMF itself, and in the bottom part we examine the relative difference between all HMFs and the DMO HMF. Grey shaded regions show deviations of $2\%$ and $5\%$. The results for the unscaled hydro masses are similar to the results presented in Schaye2023, but now presented for $M_{\mathrm{500c}}$: the different FLAMINGO simulations shift the HMF downward by up to $\approx 20\%$ due to feedback (solid lines), with the effect extending to higher halo masses for simulations with stronger AGN feedback. However, when we divide the masses by the predicted retained mass fraction for their mean halo baryon fractions (dashed lines), the effects of feedback are compensated over more than two decades of halo mass, and we recover the DMO HMF to percent-level accuracy. While the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation used here was fit only to haloes from the L2p8_m9 simulation with $M_{\mathrm{500c}}\geq 10^{12}\,\mathrm{M_{\odot}}$, it applies to all FLAMINGO simulations, and divergence at lower halo masses is slow.
Figure 17: The $M_{\mathrm{500c}}$ halo mass function (HMF) for L1_m9_DMO and various hydrodynamical variations. Shown as dashed lines are the HMFs we obtain when we divide the halo masses in the hydrodynamical simulations by their expected retained mass fractions, obtained from their baryon fractions. The bottom panel shows that without this correction (solid lines), the HMF shifts downward by up to $\approx 20\%$ depending on the strength of feedback, over a wide range of mass scales. Grey shaded regions show deviations of $2\%$ and $5\%$. After correcting the hydrodynamical masses with the retained mass fractions (dashed lines), the DMO HMF is recovered to percent-level accuracy over more than two decades of mass for all simulations.
We note that Euclid2024 recently also explored a relation to map halo masses to their DMO equivalent to recover the (cumulative) HMF, and they too start from what we call the corrected baryon fraction. However, their equivalent of the retained mass fraction is defined by adding a parameter $\delta_{f}$ to the measured baryon fraction instead, which they calibrate to simulations. The accuracy they obtain is similar to ours, however, the relation of the current work is applicable over a much wider range in halo mass scales and feedback strengths, without requiring recalibration.
@@ 4.4 Strengths and limitations
The main strength of the resummation model is that it allows halo baryon fractions to be mapped directly to a matter power suppression signal with zero free parameters, meaning that observed baryon fractions could be straightforwardly marginalised over to give a confidence interval around the predicted suppression. In this way it is similar to SP(k) (Salcido2023) – however, the resummation model does not assume any scaling law between baryon fractions and halo masses, and is much more flexible in the suppression signals it can generate, having not been trained on any simulated suppression signals.
@@
Compared to the baryonification/baryon-correction models (e.g. SchneiderTeyssier2015; Schneider2019; Arico2020; Arico2021; Schneider2025), which can also predict quantities beyond the suppression signal and can operate at the field level, the applicability of the resummation model is more narrow. However, given appropriate halo baryon fractions, the resummation model again has zero free parameters, at least at low redshift, and does not assume any functional forms for density profile shapes, whereas the default baryonification models have seven free parameters and do require these assumptions. For the same reasons, the resummation model compares favourably to halo-model based approaches (e.g. Semboloni2011; Zentner2013; Debackere2020; Pandey2025), including the popular hmcode (Mead2015; Mead2021), which contain many parameters that are trained on a limited set of DMO and hydrodynamical simulations, typically leaving at least one parameter modelling the strength of AGN feedback free. The resummation model, then, offers a great alternative to less flexible models for the suppression signal, and could be complementary to baryonification-based approaches: using its predictions for the retained mass fractions and suppression signals may help constrain the parameter space of baryonification models, allowing for more precise predictions of other quantities like halo pressure and temperature profiles.
@@
The main underlying assumption of our model is that the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation – linking the mean halo baryon fraction to the fraction of mass lost due to baryonic effects, relative to DMO – which applies to all well-resolved FLAMINGO haloes regardless of their baryonic physics, is truly universal, i.e. that it also applies to haloes in the real Universe. We have checked that haloes from the cosmo-OWLS (LeBrun2014) and BAHAMAS (McCarthy2017) simulations indeed lie on top of the same relation we present here – despite their cosmic baryon fractions lying outside the range probed by FLAMINGO– and encourage other simulators to do the same. It would be particularly interesting to see if this relation holds for simulations with different gravity and/or hydrodynamics solvers or subgrid implementations than those used here, as these may impact the result of relaxation. The results of Appendix D do indicate that the relation may shift by $\sim 1$ percentage point for lower-mass haloes when the simulation’s mass resolution is improved by a factor of 8, which we will investigate further in future work.
@@
The second-most important assumption of our model is that the non-resolved-halo matter component is sub-dominant on non-linear scales, and that we therefore do not need to accurately model the distribution of ejected matter. As we have shown in §3.6, this assumption breaks down at high redshifts, although our results there indicate that adding a possibly also universal but redshift-dependent parameter ($f_{\mathrm{ret,5x500c}}$}) could solve this issue. Some less important assumptions taken include that the linear halo bias is preserved as haloes lose mass (though we find that the relative change in bias is typically $\lesssim 1\%$), and that the halo positions do not change due to feedback in a way that significantly impacts their clustering (justified by the results of vanDaalen2014; Zennaro2025).
@@ 5 Summary
In order to derive unbiased cosmological parameters from Stage-IV weak lensing surveys like Euclid, we need models that can predict the matter power spectrum up to at least $k\,{\approx}\,{10}\,h\mathrm{\,Mpc^{-1}}$ with percent-level accuracy. The main challenge in this endeavour, as first put forward in vanDaalen2011, is that baryonic feedback significantly redistributes matter on large scales, relative to dark matter alone, the details of which are still not understood to the required level of precision. Many modelling efforts and mitigation strategies based on the results of hydrodynamical simulations exist, but these all assume functional forms for halo density profiles or the suppression signal itself, have free parameters that need to be marginalised over, or both. Here, we presented an improved version of the resummation model first described in vanLoonvanDaalen2024, which maps observed halo baryon fractions of groups and clusters to a flexible suppression signal without any free parameters, with a typical accuracy of $\lesssim 1\%$.
@@
Conceptually, the model takes observed halo masses and baryon fractions within one or more spherical overdensity radii $R_{\Delta}$, translates these into retained mass fractions $f_{\mathrm{ret},i,\Delta}\equiv M_{\mathrm{hydro},i,\Delta}/M_{\mathrm{DMO},i,\Delta}$ for different halo mass bins $i$ through the seemingly mass-independent and feedback-independent $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation presented in §2.2.1, and uses these to rescale DMO halo-matter cross spectra to account for the mass lost due to baryonic feedback. The mass removed from haloes is assumed to cluster similarly to matter already outside haloes, allowing us to rescale the DMO non-resolved-halo cross spectrum as well. All these cross spectra are then resummed to obtain a new matter power spectrum including baryons. Extensions to larger or smaller scales are straightforward (see §2.2.2-2.2.3), and allow for $\sim 1\%$ accuracy up to $k\,{\approx}\,{25}\,h\mathrm{\,Mpc^{-1}}$ given stellar and gas fractions within the inner halo. The steps of the model are summarised in §2.2.4, and a Python package called resummation is made publicly available. For a given FLAMINGO cosmology, the model maps halo baryon fractions to power suppression predictions in $1\,\mathrm{ms}$, allowing for fast iterations to marginalise over the uncertainties of observed baryon fractions of each mass bin independently.
@@
We have quantified the contributions to the suppression signal of haloes of different masses (§3.4), explaining and confirming the findings of e.g. vanDaalen2020 on the dominant role played by haloes around $M_{\mathrm{500c}}=10^{14}\,\mathrm{M_{\odot}}$, in line with vanLoonvanDaalen2024. We have shown that the baryon fraction within $R_{\mathrm{500c}}$ is also a strong predictor of the retained mass fractions on larger scales as well (§3.5), allowing the resummation model to yield accurate predictions even when only $f_{\mathrm{b,500c}}(M_{\mathrm{500c}})$ is available. We have also shown that the model’s assumption about the minor role of non-resolved-halo matter starts to break down at high redshift, but that this can be accounted for as well (§3.6) – though whether this too is model-independent remains to be seen. Finally, in Section 4, we have shown that halo relaxation following ejection likely shapes the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation (§4.1), applied our model to observed baryon fractions within $R_{\mathrm{500c}}$ to show the suppression signals they predict (§4.2), shown that the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation can also be used to accurately recover the DMO HMF from the observed halo mass function (§4.3), and discussed the strengths and limitations of the model (§4.4).
@@
Future work on the resummation model could include: improved modelling of the non-resolved-halo component, particularly for high $z$; emulation of DMO halo-matter cross spectra; investigating the dependence of the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation on resolution, numerical solvers and subgrid implementation, as well as its dependence on $\Omega_{\mathrm{b}}/\Omega_{\mathrm{m}}$ outside the limited range probed by FLAMINGO; reformulating the model in terms of multiples of $R_{\mathrm{500c}}$ and/or more readily observable halo mass definitions; inverting the model to yield baryon fractions given a suppression signal, without degeneracies; and improving the model in the presence of asymmetric redistributions of matter, such as through jets.
@@ Acknowledgements
We thank Jaime Salcido for providing the baryon fractions and Monte Carlo errors derived from Akino2022, converted to the D3A cosmology. We thank Ajay Dev for providing the baryon fractions and percentiles derived from the combined dataset of Dev2024 and a useful discussion on their cosmological and mass definition dependencies. We also thank the anonymous referee for their useful comments, which have led to the improvement of this manuscript.
@@
The halo-matter DMO cross spectra, as well as baryon fractions and stellar fractions for the FLAMINGO simulations used in this work, are publicly available as part of the Python package resummation. This can be installed via the PyPI package index (pip install resummation) or downloaded from GitHub [^15] https://github.com/FLAMINGOSIM/Resummation, which includes examples. The matter power spectra of all FLAMINGO simulations are available on the FLAMINGO project’s web page. [^16] https://flamingo.strw.leidenuniv.nl/
@@ Appendix C Fitted results at higher redshifts
That there exists a redshift dependence at all, hints at other dependencies of the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation. Elbers2025 show that the retained mass fraction, there indicated as $R_{\mathrm{b}}$, depends on halo concentration and formation time at fixed DMO mass. While the relation presented in this work likely absorbs some of those dependencies, some residual dependence probably remains.
@@ Appendix D Sensitivity to changes in resolution
The higher-resolution L1_m8 simulation also agrees for massive haloes. However, as the halo mass decreases it starts to undershoot the m9 simulations by $1$ percentage point consistently for $R_{\mathrm{200m}}$ regions. For $R_{\mathrm{500c}}$, the deviation is larger, though it still does not exceed $\approx 2$ percentage points at any halo mass. A similar discrepancy between mass ratios in these FLAMINGO simulations was recently noted by Schneider2025.
Figure 21: The result of applying the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation fit to L2p8_m9 to the higher-resolution L1_m8 simulation. Left: The true suppression signal compared to the model prediction. Right: The ratio of the model to true suppression signal. As expected from the results of Fig. 20, the universal relation overpredicts the retained mass fraction for this resolution, although the prediction is still $\lesssim 1\%$ accurate for $k\,{\lesssim}\,{3}\,h\mathrm{\,Mpc^{-1}}$.
@@
We compare the model prediction for the suppression signal for L1_m8 to the true one in Fig. 21, using the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation fit to L2p8_m9. The model tends to underestimate the amount of suppression, as expected from the results of Fig. 20, since the retained mass fraction is overestimated. Still, the prediction is $\lesssim 1\%$ accurate for $k\,{\lesssim}\,{3}\,h\mathrm{\,Mpc^{-1}}$, and stays below $4\%$ for all $k\,{\leq}\,{10}\,h\mathrm{\,Mpc^{-1}}$.
We note that the resolution is not the only property of the simulation that changes between L1_m9 and L1_m8: as noted in Schaye2023, two changes were made to the simulation code after most of the FLAMINGO simulations were completed. First, there was an update to the simulation code that fixed the repositioning of supermassive black holes, to keep them in the centre of their halo’s potential. In most FLAMINGO simulations, the black hole’s own potential was erroneously not excluded from the potential calculation (see their §2.3.5). Second, in these same simulations, particles with exactly zero metallicity had their star formation suppressed, although this only significantly affected the very lowest-mass galaxies (see their footnote 5). Of the simulations explored here, only L1_m8, L1_m10 and the jet simulations have the black hole potential correctly subtracted, and the star formation bug fixed. In our results, these are among the most divergent simulations in terms of model performance, which the code changes (particularly the black hole repositioning) could possibly have contributed to.